OpenStudy (anonymous):

d=54ft,t=27 seconds r= ft/sec is?

OpenStudy (anonymous):

54/27

OpenStudy (stamp):

d=rt r=d/t r = ft/sec? ft = 54 sec = 27 r = 54/27 ft /sec